OpenStudy (anonymous):

The functions f, g and h if graphs are drawn in the coordinate below the regulation of the form p (x) = ax ^ 2 + bx + c Determine the sign of a, b​​, c and d (where d is the discriminant). Give reasons for your answer.

OpenStudy (anonymous):

This is the graph

OpenStudy (anonymous):

@whpalmer4, is this something you can help me with?

OpenStudy (anonymous):

or you, @AravindG ?

OpenStudy (aravindg):

For all cup shaped quadratics a is positive. When it is inverted sign of a is negative.

OpenStudy (anonymous):

Okay, that makes sense... What about b, c and d? @AravindG

OpenStudy (aravindg):

Let's deal with d first. When graph cuts the x axis at two points we say 2 roots occur and hence d>0. When it cuts in only 1 point d=0 and when it doesnt cut any point d<0.

OpenStudy (aravindg):

Keeping in mind that d=b^2-4ac and the sign of a, you can predict sign of b,c for the functions given.

OpenStudy (anonymous):

So for an example, graph g: a is negative and d=0?

OpenStudy (anonymous):

No, wait d<0 for d, because it doesn't cut the x axis, right? @AravindG

OpenStudy (aravindg):

^^Exactly. You get a medal for that :)

OpenStudy (anonymous):

Haha, thanks! it's the same for f too? It doesn't look like it's going the x axis, and a should be positive, right? And graph h, a is positive with d>0.. right? @AravindG How do you find b then? isn't c, where the graph cuts the y-axis?

OpenStudy (aravindg):

For f it definitely touches x axis at one point. So it is not same cases as g. Yes a is positive for f.

OpenStudy (aravindg):

You are right about h by the way :) ok I will show you an example of finding sign of b and c taking the function f: d=b^2-4ac d=0 as cuts only in 1 point. This means b^2=4ac Now we know a is positive and b^2 is always positive. Hence c is also positive.

OpenStudy (whpalmer4):

I'm awarding a medal for the same reason, but I must say that @AravindG 's write-up here is very fine, and I wish I could award two medals!

Parth (parthkohli):

(If I may add another method). Let's find the sign of \(b\) in \(h(x)\). Both roots positive, so the sum of roots is positive. So \(-b/a\) is positive. You know the sign of \(a\) (positive), so \(b\) must be negative.

Parth (parthkohli):

Similarly the sum of roots in f is negative...

OpenStudy (anonymous):

@AravindG, I'm not fully with you... How did you get from d=b^2-4ac to b^2=4ac...

OpenStudy (anonymous):

@ParthKohli, so what you're saying is that if a is positive b should be negative? always?

OpenStudy (aravindg):

Thank you @whpalmer4 :) Recall that for a graph that cuts only at one point d=0 Put it in equation for d ie d=b^2-4ac-->b^2=4ac

Parth (parthkohli):

\[-b/a > 0\]-b/a is the sum of roots; a is positive because the graph points up so we can just remove the positive denom from the inequality\[-b> 0\]\[b < 0\]this is for h(x).

Parth (parthkohli):

do you see how this was done?

OpenStudy (anonymous):

A little confused.. @ParthKohli

Parth (parthkohli):

hmm, do you know the formula sum of roots and product of roots of a quadratic equation?

Parth (parthkohli):

sum of roots is -b/a product of roots is c/a ^ that

OpenStudy (anonymous):

@AravindG so if i do that, i'll know a,b,c and d?

OpenStudy (anonymous):

No, but I guess I know now? @ParthKohli -

Parth (parthkohli):

but this is only an alternative. so you should follow what you were taught.

OpenStudy (anonymous):

@ParthKohli, I take online lessons - and there isn't really much of an method :/

OpenStudy (aravindg):

You get signs of a,c,d. For sign of b I just look at the axis of symmetry. If it is having a negative x value b becomes positive and vice versa.

OpenStudy (aravindg):

^For a<0 Do the opposite of what is said above.

OpenStudy (anonymous):

@AravindG so if a was positive, b would be negative? But how do I explain how i know this?

Parth (parthkohli):

no sweat :) it's a very useful result though. here's how we derive it: let \(p\) and \(q\) be the roots/zeroes of a quadratic \(x^2 + bx + c= a\left(x^2 + b/a ~x ~+ c/a\right)\). by the factor theorem, the polynomial can be written as \(a(x-p)(x-q) = a(x^2 - (p+q)x + pq)\). so \(a(x^2 + b/a ~x + c/a) =a(x^2 - (p+q)x + pq)\). look at the coefficients and see how\[p+q = b/a\\pq = c/a\]you can use this now to show what aravindg said.

Parth (parthkohli):

you can also use this to find c!

Parth (parthkohli):

but c is the value of y when x = 0, so let's keep it simple lol

OpenStudy (aravindg):

Using the axis of symmetry concept! Always draw a vertical line that cuts the graph into two. Look at the point where it touches x axis. The sign of this value is important. For example For graph h it is negative. Once you know the sign its easy x intercept of axis of symmetry=-b/2a Put in the sign and we can do all comparisons. For h, -ve=-b/2(+ve) Hence b=-ve.

OpenStudy (anonymous):

@AravindG, okay! Thanks! I'll see if i can write into words in my language, lol. @Parthkohli Thanks a bunch too! I'll tag you if i need more help, lol.

OpenStudy (aravindg):

yw :) Indeed these notes can prove very useful in the end exam. Good luck!

OpenStudy (anonymous):

@AravindG, yes, indeed! Thanks!